a body of mass 2 kg is floating in water with half of its volume outside the surface of water how much force is required to just submerge the cube in water...ans 20N......plzz explain how ..?

Grade:11

2 Answers

Mahima Kanawat
5 years ago
Dear student 
Buoyancy force is acting upward on the body by the fluid which makes the body float.
To submerge it the force equal to buoyancy force shoulde be applied on opposite direction.
Buoyancy force = mg = 2*10 =20 N
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT 
Anupam Mahato
2 years ago
We know buoyant force= weight of water displaced
Weight of body= 2kg × 9.8m/s²= 19.6N= weight of water displaced
 
So buoyant force= 19.6N